OpenStudy (anonymous):

Subtract/simplify completely: 1/3y-2/4x

OpenStudy (anonymous):

1. Problem: Solve the following system: x + y = 11 3x - y = 5 Solution: Solve the first equation for y (you could solve for x - it doesn't matter). y = 11 - x Now, substitute 11 - x for y in the second equation. This gives the equation one variable, which earlier algebra work has taught you how to do. 3x - (11 - x) = 5 3x - 11 + x = 5 4x = 16 x = 4 Now, substitute 4 for x in either equation and solve for y. (We use the first equation below.) 4 + y = 11 y = 7 The solution is the ordered pair, (4, 7).

OpenStudy (anonymous):

\[\frac {1} {3y} - \frac {2} {4x} = \frac {4x} {12xy} - \frac {6y} {12xy} = \frac {4x - 6y} {12xy} \]

OpenStudy (anonymous):

You're great! Thank you!

OpenStudy (anonymous):

Actually it simplifies to: \[\frac {2(2x- 3)} {2(6xy)} = \frac {2x-3} {6xy} \]

OpenStudy (anonymous):

Is there be a missing y following the 3?

OpenStudy (anonymous):

Hmmm, yeah, forgot the y